THE A^{1}Σ^{+} − X^{1}Σ^{+} AND A^{′ 1}Π − X^{1}Σ^{+} CHEMILUMINESCENCE SPECTRA OF SrO FROM A FOURIER TRANSFORM SPECTROMETER

\underline{HONGZHI LI}, RANDALL SKELTON, CRISTIAN FOCSA, BERNARD

PINCHEMEL, and PETER F. BERNATH, Guelph-Waterloo Centre for Graduate Work in Chemistry and Biochemistry, Department of Chemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada.

The A^{1}Σ^{+} − X^{1}Σ^{+} and A^{′ 1}Π − X^{1}Σ^{+} spectra of SrO were observed in the range of 3,900 to 13,600 cm^{−1}at a resolution of about 0.03 cm^{−1}. The SrO chemiluminescence was produced in a Broida-type oven from the Sr + N_{2}O reaction and detected with a Fourier transform spectrometer. A total of 75 bands from ^{88}SrO, ^{87}SrO and ^{86}SrO were measured for the A-X transition, and vibrational levels of the ground state were analyzed up to v′′ = 12. Over 10,000 rovibrational lines from the A-X transition were fitted to obtain significantly inproved spectral constants of the ground state. Strong perturbations were observed in the A state. The A′-X infrared emission spectrum was also analyzed. A set of Dunham parameters for the A′ state was obtained through the fitting of the data.